A bag contains 30 red tiles, 15 green tiles, and 5 yellow tiles. One tile is drawn and then replaced. Then a second tile is drawn. What is the probability that the first tile is yellow and the second tile is green?
Answer by stanbon(75887)

A bag contains 30 red tiles, 15 green tiles, and 5 yellow tiles. One tile is drawn and then replaced. Then a second tile is drawn. What is the probability that the first tile is yellow and the second tile is green?
----
Because of replacement, these events are independent.
-----
P(yellow AND green) = P(yellow)*P(green) = (5/50)(15/50) = (1/10)(3/10) = 3/100
----------------------------------
Cheers,
Stan H.
====================